1-) At December 31, 2016, Blossom Company Imports reported this information on its balance sheet.

Accounts receivable $581,200
Less: Allowance for doubtful accounts 38,440


During 2017, the company had the following transactions related to receivables.

1. Sales on account $2,863,600
2. Sales returns and allowances 72,240
3. Collections of accounts receivable 2,512,000
4. Write-offs of accounts receivable deemed uncollectible 42,630
5. Recovery of bad debts previously written off as uncollectible 14,180

Compute the accounts receivable turnover=???. (Round answer to 1 decimal place, e.g. 12.5.)

Compute the average collection period= ??? (Round answer to 1 decimal place, e.g. 12.5.)

2-)

June 1 Purchased books on account for $1,065 (including freight) from Catlin Publishers, terms 2/10, n/30.
3 Sold books on account to Garfunkel Bookstore for $1,900. The cost of the merchandise sold was $700.
6 Received $65 credit for books returned to Catlin Publishers.
9 Paid Catlin Publishers in full.
15 Received payment in full from Garfunkel Bookstore.
17 Sold books on account to Bell Tower for $1,900. The cost of the merchandise sold was $750.
20 Purchased books on account for $700 from Priceless Book Publishers, terms 1/15, n/30.
24 Received payment in full from Bell Tower.
26 Paid Priceless Book Publishers in full.
28 Sold books on account to General Bookstore for $1,250. The cost of the merchandise sold was $810.
30 Granted General Bookstore $270 credit for books returned costing $65.

June -9

Account payable 1000
inventory ?
cash ?

june - 24

cash ?
sales discounts ?
acct receivable 1900

june -26

account payable 700
inventory ?
cash ?
Answer #1

1.
Accounts Receivable ending balance = $581200+2863600-72240-2512000-42630+14180-14180 = $817930
Average accounts receivable = ($581200+817930)/2 = $699565

Accounts Receivable turnover = Net Sales / Average Accounts Receivable
= (2863600-72240) / 699565 = 4.0

Average Collection Period = 365/4 = 91.3 days

2.

No discount on sales given

Accounts Payable 1000
       Inventory 20
       Cash 980
Cash 1900
Sales Discount 0
     Accounts Receivable 1900
Accounts Payable 700
       Inventory 7
       Cash 693